FEL Upgrade Project Weekly Brief

December 2-6, 2002

Installation, check-out and commissioning activities continued through a very busy week:

Highlights include: (1) vacuum qualification of the 180 bend chamber for Arc 1; (2) successful high voltage from the upgraded HPA that feeds the injector cryounit which readies us for SRF commissioning of the cryounit next week (3) successful check-out of the cathode alignment and low current activation of the cesiator that allows the gun bakeout process to proceed; (4) the scanner for imaging the quantum efficiency of the GaAs photocathode has been upgraded and re-commissioned.

In preparation for our biweekly installation coordination meeting this week, we queried both our external and internal suppliers for updated completion dates for the qualification and installation of the remaining hardware for the Upgrade. This information was also forwarded to DOE and ONR. We continue to try to optimize our use of the available labor and resources to ensure that all necessary hardware is in place for first the lasing attempts before the CEBAF shutdown begins on Feb.10. (When vacuum and alignment labor become unavailable for 6-8 weeks).

We prepared and sent to the DOE Contracting Officer and the ONR Program Office a requested estimate of the close-out costs for the current FY02-03 Navy FEL program. We also forwarded a response to a series of questions from the ONR Program Office concerning the status of the magnet/vacuum chamber qualification/installation process, and the photoinjector specifications.

Two white paper proposals were prepared and submitted this week to the DoD Joint Technology Office (JTO) on the topics of (1) "FEL Higher Order Mode and Beam Breakup Model Validation" and (2) "Producing Damage-Resistant NIR Optics for High Average Power Ultra-Short-Pulse Free Electron Lasers".

George Neil and Michelle Shinn are scheduled to give presentations on the results of our two FY02 JTO projects at the JTO Program Office in Albuquerque next Tuesday (Dec.10) on (1) cryo-cooled mirrors for high average power applications; and (2) development of an advanced drive laser for high power photoinjectors. Key test data on the cryo-cooled concept is being obtained today to keep up the group’s penchant for "just-in-time" delivery.

WBS 4 (Injector):
The cathode stalk was run in and seated in position. An optical check was made through the lightbox and everything looks good for both the lightbox and the wafer. Characterized the low end degas cycle of the cesium channels. While in the upper end, heating caused large pressure

transients in the unbaked gun which precluded us from checking for a photo response prior to a bake. The gun and downstream lightbox regions are being set up for vacuum bakes.

Gun HVPS - The Conditioning/Running Resistor assembly is completed and ready for operations. The final assembly and test of the HVPS with its EPICS interface box is scheduled for next week during the Gun bake period.

WBS 6 (RF):
Quarter HVPS - Completed crowbar testing for unit #4. A loose ground connection was found in unit #4 and protective circuits were added to both units. Completed taking Miram curve data for both units. The data will be plotted and new operating points will be selected. Both units will be producing RF into shorts early next week. The SRF checkout will be able to start next Wednesday evening. There is still some debugging to be done with the software, but progress is being made.

WBS 8 (Instrumentation):
The devil is in the details. Or, in our case the angel, many of the subsystems are getting signed off as operational. The latest was the upgraded photo cathode scanner. This has the new log amp PC board that had been developed for the IR demo but never found it's way into service. The VME hardware for this is also common to the M56/bunch length monitor system. High level software checkout is in progress for both of these systems.

The dump water skid with the new conductivity meters is also operational. These units control the polishing loop to maintain the water at ~2 Meg-ohm cm. The instruments that monitor the individual dumps for the MPS are ready and awaiting the water connections.

DC power leads for the 2nd Arc string were installed up to the 166kW box supply. The Thermal Interlock chassis for the 1st Arc string has been checked out successfully. High level check out of the 5kW Dispersion power supply and associated screens was successful.

Installation of the magnet shunt cable for the 1st, 2nd Arc strings and optical chicane has been completed, termination's into the 1st Arc string shunt chassis should begin soon. We are assisting in the installation of an additional 66kW box supply in the test lab magnet test area to help prevent a bottle-neck in measurements come January.

Low level checkout and alignment of the wiggler viewers is complete. The wiring diagram for the BLM system has been signed off. Schematic capture for the stepper motor front panels boards were sent out to the vendor with a 3 day turn. The first of the new stepper motor chassis is in operation without the new panel for system checks. The motor cables are complete for the high reflector (HR) end of the optical cavity and will be checked out when cavity is complete. An LVDT cable for the HR was found to have insufficient shielding, new fully individually shielded cable is due in 12/6 and will be installed and tested early next week. The rest of the VME system for monitoring the position for the HR & OC mirrors of the optical cavity has been tested with local LVDTs.

Machine Protection System (MPS) - The final boards for fiber optic inputs and outputs are in, fabricated, and being tested. The complete specification of the protection matrix is nearly complete. There are a couple miscellaneous input such as dump water, raster power supply, and

laser safety system that remain to have channel assignments and cable definitions made - this will be completed next week.

Beam Loss Monitors (BLMs) - EES has provided us the second BLM board. It is being installed and tested. There are two more coming plus a spare. The cabling in the tunnel is complete, the heads are fragile and will be installed as needed for commissioning. All of the heads are complete and calibrated.

Drive Laser Pulse Controller Upgrade (DLPC) - The RF cables for the micropulse generator board (the VME CO-305) are being made, when completed, the board will be installed into IOCFEL10 Also, the ioc(10) is being fitted with a -5vdc power supply as is needed for the new card.

WBS 9 (Transport):
Work Arounds
Substitution of IR Demo 180° Bend (DY) for the Upgrade 180° Bend (GY) in the First Arc.
• The chamber from the DY is qualified by the vacuum group and will be inserted back into the
DY next week.
• The stands for the DY are being installed in the enclosure. The stands are inboard of the wider stance of the stands for the final GY magnet. Those stands will not be disturbed.
Diversion of the beam after the Wiggler.
• The vacuum group is working on the vacuum pipes and the stand is in.
• The original IR Demo, DW chamber has been modified for this application.
• The GW magnet, planned to go into this position is in transit from Oakland and will be put in position without measurement.
Optical Chicane Dipoles (GW)
• The 6 GW Dipoles for the UV machine shipped from Oakland on Monday, December 2 and should arrive late Friday or on Monday.
Injector Dipoles (DU/DV)
• The Jlab Shop will take on the task of removing and re gluing the Purcell Gap shims on the second GU dipole. The bond has been disturbed (perhaps during machining of the edges to shorten the magnet (to match field integral to core field) and the slight bubbles produce field that is out of specification. They will have it ready by the time it is needed for recirculation.
Arc 180 Degree Dipoles (GY)
• Wang NMR impregnated the coil with epoxy last week, but they were on the raw edge of not having enough epoxy. It initially filled the pump out sprews in the mold, but over night that level fell below the exterior of the sprew tube. They are getting more epoxy and plan to complete impregnation and cure by early next week. There is no danger of the epoxy prematurely setting up because it is a premixed, heat-activated epoxy. This delay and the learning curve in getting the coil into the mold have set their ship date to December 19.
• Wang NMR is gluing the Purcell gap shims on top yoke.
Arc Bend, Reverse Bend Dipoles (GQ, GX)
• The first GQ remains on the test stand. It passed core field specification at 145 MeV. However, Field integral exhibits a Sextupole of up to 4 parts per ten thousand. And the 80 MeV integral is 5.5. We are working on a three-pronged attack. 1. We tried setting current from the bottom of the hysteresis loop. All flatness at all excitations ranged between 2 and 4 parts per ten thousand ad the core fields reversed direction of their error from a "frown" to a "smile". It is not the best fix because the value of the integral is out of range of the field clamp adjustment and varies with excitation. 2. We added shims to the sides of the end faces of the poles and as a more practical alternate to the center of the edges of the field clamps. Both systems work to flatten integrals within spec and we prefer the later. 3. However core fields are still twice spec at 80 MeV/c. We will try setting the current in dithering method, decreasing the overshoot either side of the set value for several cycles much like CEBAF is setting some Quadrupoles. This method may set the core fields to flat. Stay tuned.
• Wang is ready to impregnate the fourth GX coil once they get epoxy and is preparing the mold for the GQ coils.
Trim Quad (QT)
• The trim coil to create the GC vertical corrector function within the trim quad is being sent out for bid.
Sextupole (SF)
• No progress
Octupole (OT)
• We are extracting coils from the BT correctors for use on these magnets.
Beam Line and Vacuum
• Chicane chamber fabrication in the Jlab Shop continued

WBS 11 (Optics):
Work this week concentrated on establishing utilities (electrical and cooling) to the optical cavity internals. Over the holiday we vacuum-tested the connectors to ensure they were not impregnated with hydrocarbons, and confirmed they were not. Work is concentrating on the HR assembly.

The fabrication of parts for the diagnostic viewer/laser safety shutter and the insertable mirrors are being coordinated by our Machine Shop. Deliver before Christmas appears tight, so we may have to delay installation to the first of the year. This shouldn't impact the lasing schedule.

We plan on doing a test of our cryocooled mirror today (12/06) This is an important milestone for our JTO-sponsored efforts as well as a required component for the UV FEL. We worked on close-out presentations to give to the JTO.